OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation x/2+1/4=3/4

OpenStudy (bradely):

x/2+1/4=3/4 Subtract 1/4 from each side x/2+1/4-1/4=3/4-1/4 x/2 =3/4-1/4 x/2 =(3-1)/4 x/2 =2/4 x =1 Source: http://www.mathskey.com/question2answer/

OpenStudy (mathstudent55):

The denominators are 2, 4, and 4. What is the smallest number that you can divide evenly by 2 and 4?

OpenStudy (anonymous):

\[\frac{ x }{ 2 } + \frac{ 1 }{ 4 } = \frac{ 3 }{ 4 }\]

OpenStudy (mathstudent55):

@bradely If this were a question on a test, you'd get no credit for it because even though you solved the equation correctly, you didn't answer the question being asked.

OpenStudy (anonymous):

I already tried to put in 1 but it says its wrong. I don't understand how I'm getting it wrong.

OpenStudy (mathstudent55):

Here is an example. Let's use the number 16 as a common denominator. 16/2 = 8 (no remainder) 16/4 = 4 (no remainder) 16 is a common denominator for 2 and 4 because when you divide 16 by 2 or by 4 you get no remainder. 16 is not the least common denominator. This problem is asking for the least common denominator (LCD) of 2 and 4. What is the smallest number that you can divide by 2 and divide by 4 and get no remainder? That is what an LCD is.

OpenStudy (anonymous):

2, 4, and 4. Which is the smallest number?

OpenStudy (mathstudent55):

One way to find the least common denominator of numbers is to list the multiples of the numbers. Take each number, and multiply it by 2, 3, 4, 5, 6, ... Write those multiples. Then look for the smallest multiple that is common to both.

OpenStudy (mathstudent55):

Here is an example. Find the LCD of 3 and 5. Solution Multiples of 3: 3, 6, 9, 12, 15, 18, 21, ... Multiples of 5: 5, 10, 15, 20, 25, ... Smallest multiple of both 3 and 5: 15 15 is the LCD of 3 and 5 Now do the same for 2 and 4.

OpenStudy (triciaal):

@kspencer you did not answer the question asked. read the question again then follow what @mathstudent55 has above